OpenStudy (anonymous):

Simplify. (2x − 5)(6x + 5) + 13

OpenStudy (anonymous):

(2x − 5)(6x + 5) + 13 = 2x(6x + 5) -5(6x + 5) + 13 = 12x^2 + 10x - 30x - 25 + 13 = 12x^2 - 20x - 12